# Version of Tuesday $24^{\text{th}}$ June, 2025

### v1

@@ Abstract.
In this work, we solve a discrete optimal transport problem in a nonuniform environment. The key challenge is to form the cost matrix, which requires finding the optimal path between two points, and for this task we formulate and solve the associated Euler-Lagrange equations. A main theoretical result of ours is to provide verifiable sufficient conditions of optimality of the solution of the Euler-Lagrange equation. We propose new algorithms to solve the problem, and illustrate our results and performance of the algorithms on several numerical examples.
@@
Optimal transport, optimal trajectory, calculus of variations, necessary and sufficient optimality condition, assignment problem, Sinkhorn method
@@ 1. Introduction
Motivated by solving optimal transport problems in the presence of obstacles, in this work we consider the discrete optimal transport of masses in a nonuniform environment.
@@
As with all optimal transport problems, a main issue is what is the transportation cost $c({\mathbf{x}},{\mathbf{y}})$ of moving one unit of mass from location $\mathbf{x}$} to location $\mathbf{y}$}, where $\mathbf{x}$} and $\mathbf{y}$} are points in some subset of ${\mathbb{R}}^{n}$. In a uniform environment, the case that is usually considered in the literature, the transportation cost $c({\mathbf{x}},{\mathbf{y}})$ is typically given by a $p$-norm, $c({\mathbf{x}},{\mathbf{y}})=\|{\mathbf{x}}-{\mathbf{y}}\|_{p}$, $1<p<\infty$, the $2$-norm $\|{\mathbf{x}}-{\mathbf{y}}\|_{2}$, or simply $\|{\mathbf{x}}-{\mathbf{y}}\|$, being the most obvious choice, and also by the $2$-norm square, $c({\mathbf{x}},{\mathbf{y}})=\frac{1}{2}\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}$, a case that is known as the $2$-Wasserstein distance, or simply Wasserstein distance. In a uniform environment in ${\mathbb{R}}^{n}$}, computation of the cost is straightforward. But, in a nonuniform environment, the cost will generally depend on the path followed to go from $\mathbf{x}$} to $\mathbf{y}$}, and this makes the computation of $c({\mathbf{x}},{\mathbf{y}})$ a challenge in its own rights.
@@
Optimal transport, in one of its several variants, has been receiving a lot of attention in recent decades, in no small part because of its flexibility to adapt to many problems of seemingly different nature, such as the Schrödinger bridge problem, unbalanced optimal transport, transport in the presence of physical constraints, and the use of optimal transport in Machine Learning (e.g., see [15]). Our work is particularly motivated by studying optimal transport in the presence of physical constraints, obstacles. The study of optimal transport in the presence of obstacles has received both theoretical and computational attention; for example, see [4] for theoretical results, and [1, 3] for computational techniques. In particular, Barrett and Prigozhin are concerned with transporting a mass into another, in the plane, by minimizing the cost given by the Euclidean distance when there is an obstacle along straight lines paths, and their computational approach consists of a regularization of the cost, followed by a finite element discretization of the underlying Monge-Ampere PDE. Presence of an obstacle is one instance of a nonuniform environment, which is the general case we consider, and we do it in the context of a discrete optimal transport problem so that it becomes necessary to compute optimal trajectories between two points (in the plane or in space) subject to a weighted “distance” function. This very task bears similarities with that of obtaining geodesics on general Riemannian manifolds (e.g., see [9, Section 2.C]), as we will exemplify in the body of the paper.
@@
A plan of the paper is as follows. In Section 2, we review the basic setup of the discrete optimal transport problem and discuss the two choices of cost we consider: the length and the energy costs, generalizing the cases of the Euclidean distance and the $2$-Wasserstein distance to nonuniform environments. In Section 3, we derive the Euler-Lagrange equations for the cases of interest to us, recall a well known equivalence result between the two cases considered, and discuss its implications for algorithmic development. We also give an important new theoretical result establishing sufficient conditions for optimality. In Section 4, we outline the algorithms adopted in full details, and in Section 5 we present simulation results on several examples, highlighting performance on a number of scenarios and challenges that must be overcome.
2. Setup of the problem and our tasks
@@
2.1. Discrete Optimal Transport
This is the most commonly studied setup of an optimal transport problem. In no small part because of widely available, and fairly robust, numerical methods for its solution.
@@ 2.1. Discrete Optimal Transport
2.2. Cost Matrix
In general, setting up the cost matrix is the most difficult task as we discuss below. First of all, we need to select a cost $c({\mathbf{a}},{\mathbf{b}})$, which is the cost to move one unit of mass from location ${\mathbf{a}}\in{{\mathbb{R}}^{n}}$ to location ${\mathbf{b}}\in{{\mathbb{R}}^{n}}$, accounting for the environment. We consider two costs, the “length”, and the “energy” costs, which generalize the Euclidean and the Wasserstein distances, respectively.
2.2.1. The length cost
This is given by
@@ 2.2.1. The length cost
where the norm is the Euclidean norm. The kernel function $K$ is what characterizes the environment, that is its geometry. If $K=1$ (or any fixed constant, for what matters), we have a uniform environment, otherwise a nonuniform environment. Quite clearly, if $K=1$, we are just looking at the arc-length of the path from $\mathbf{a}$} to $\mathbf{b}$}, whereas if $K$ is not constant we have a weighted arc-length of the path. It is a simple exercise in calculus of variations (e.g., see [13]) that, when $K=1$, the shortest path is a straight line: ${\mathbf{x}}(t)={\mathbf{a}}+t({\mathbf{b}}-{\mathbf{a}})$, $0\leq t\leq 1$, so that the cost is precisely the Euclidean distance from $\mathbf{a}$} to $\mathbf{b}$}: $c({\mathbf{a}},{\mathbf{b}})=\|{\mathbf{a}}-{\mathbf{b}}\|$, as one should expect in a uniform environment.
2.2.2. The energy cost
This is given by
@@ 3. Optimal Trajectory
Our main concern in this section is to solve the minimization problem for the optimal trajectory from $\mathbf{a}$} to $\mathbf{b}$} for $\mathbf{x}$}-dependent weight function $K({\mathbf{x}})$, which we need to do in order to setup the cost matrix. To present the techniques we adopted, we first unify the notation.
Let ${\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})$ be the integrand in either (4) or (5); $\mathcal{L}$} is called the Lagrangian. Our task is to find the trajectory ${\mathbf{x}}\in{\mathcal{P}}({\mathbf{a}},{\mathbf{b}})$ which gives the minimum value below
@@
Regardless of whether $\mathcal{L}$} is given by (4) or by (5), the minimizing path will be called a geodesic; this naming will be justified by Fact 3.3 below.
@@ 3.1. Necessary condition: Euler-Lagrange equations
The problem formulated in (6) is a classical problem of Calculus of Variations and the search for a minimizer follows the approach based on solving the Euler-Lagrange equations, which are first order necessary conditions for obtaining a minimum. For (6), the Euler-Lagrange equations are:
(7) $\frac{\partial{\mathcal{L}}}{\partial x_{i}}-\frac{d}{dt}\bigg[\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]=0\ ,\ i=1,2,\dots,n\ .$
Multiplying (7) by $\dot{x}_{i}$, we get:
$\frac{\partial{\mathcal{L}}}{\partial x_{i}}\dot{x}_{i}-\frac{d}{dt}\bigg[\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]\dot{x}_{i}=0\implies\frac{\partial{\mathcal{L}}}{\partial x_{i}}\dot{x}_{i}=\frac{d}{dt}\bigg[\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]\dot{x}_{i}\ ,\ i=1,2,\dots,n\ .\$
Since the Lagrangian ${\mathcal{L}}({\mathbf{x}}(t),\dot{\mathbf{x}}(t))$ does not depend on $t$ explicitly, we get
$\frac{d{\mathcal{L}}}{dt}=\sum_{i=1}^{n}\bigg[\frac{\partial{\mathcal{L}}}{\partial x_{i}}\dot{x}_{i}+\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\ddot{x}_{i}\bigg]\ \quad\text{or}\quad\frac{d{\mathcal{L}}}{dt}=\sum_{i=1}^{n}\bigg[\frac{d}{dt}\bigg[\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]\dot{x}_{i}+\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\ddot{x}_{i}\bigg]\ .$
Finally, using the product rule we get the so-called first integral formula:
(8) $\frac{d{\mathcal{L}}}{dt}=\sum_{i=1}^{n}\bigg[\frac{d}{dt}\bigg(\dot{x}_{i}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg)\bigg]=\frac{d}{dt}\sum_{i=1}^{n}\dot{x}_{i}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\implies\frac{d}{dt}\bigg[\sum_{i=1}^{n}\dot{x}_{i}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}-{\mathcal{L}}\bigg]=0\ .$
The formula in (8) is conducive to obtaining a system of ODEs, as follows. Expand (8)
$\sum_{i=1}^{n}\ddot{x}_{i}{\mathcal{L}}_{\dot{x}_{i}}+\sum_{i=1}^{n}\dot{x}_{i}\bigg[\frac{d}{dt}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]-\sum_{i=1}^{n}{\mathcal{L}}_{x_{i}}\dot{x}_{i}-\sum_{i=1}^{n}{\mathcal{L}}_{\dot{x}_{i}}\ddot{x}_{i}=0\ ,$
${\text{so that we get}}\quad\sum_{i=1}^{n}\bigg[\frac{d}{dt}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}-{\mathcal{L}}_{x_{i}}\bigg]\dot{x}_{i}=0\ .$
Therefore, (8), is satisfied if (a sufficient condition) the terms in brackets vanish, that is if $\frac{d}{dt}\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}-{\mathcal{L}}_{x_{i}}=0$, for all $i$, or $\sum_{j=1}^{n}{\mathcal{L}}_{\dot{x}_{i}\dot{x}_{j}}\ddot{x}_{j}+\sum_{j=1}^{n}{\mathcal{L}}_{\dot{x}_{i}x_{j}}\dot{x}_{j}-{\mathcal{L}}_{x_{i}}=0$, for all $i$, that is
@@
where $D^{2}_{\dot{\mathbf{x}}}{\mathcal{L}}\in{{\mathbb{R}}^{n\times n}}$, $(D^{2}_{\dot{\mathbf{x}}}{\mathcal{L}})_{ij}={\mathcal{L}}_{\dot{x}_{i}\dot{x}_{j}}$, and $D_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathcal{L}}\in\mathbb{R}^{n\times n}$, $(D_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathcal{L}})_{ij}={\mathcal{L}}_{\dot{x}_{i}x_{j}}$. In short, the Euler-Lagrange equations (7) are satisfied along a path ${\mathbf{x}}\in{\mathcal{P}}({\mathbf{a}},{\mathbf{b}})$ if (9) is. To further simplify (9), we now look at the specific form of the Lagrangian $\mathcal{L}$} in the two cases of interest to us.
@@ 3.1.2. Euler-Lagrange for energy cost
Here,
(13) ${\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})=\frac{1}{2}K^{2}({\mathbf{x}})\|\dot{\mathbf{x}}\|^{2}\ .$
In this case, it is easier to rewrite the boundary value problem (9).
@@
$K^{2}({\mathbf{x}})\ddot{\mathbf{x}}+2K({\mathbf{x}})\dot{\mathbf{x}}[\nabla K]^{T}\dot{\mathbf{x}}=K({\mathbf{x}})\|\dot{\mathbf{x}}\|^{2}\nabla K$
@@
$\ddot{\mathbf{x}}=\frac{\|\dot{\mathbf{x}}\|^{2}}{K({\mathbf{x}})}\nabla K-\frac{2(\nabla K)^{T}\dot{\mathbf{x}}}{K({\mathbf{x}})}\dot{\mathbf{x}}\$
@@
An interesting and useful criterion about solutions of (14) is that they have constant speed.
@@
$\ddot{\mathbf{y}}=\frac{\|\dot{\mathbf{y}}\|^{2}}{K({\mathbf{y}})}\nabla K({\mathbf{y}})-\frac{2[\nabla K({\mathbf{y}})]^{T}\dot{\mathbf{y}}}{K({\mathbf{y}})}\dot{\mathbf{y}}$
@@
$K\ddot{\mathbf{y}}^{T}\dot{\mathbf{y}}=-\|\dot{\mathbf{y}}\|^{2}[\nabla K]^{T}\dot{\mathbf{y}}\quad\text{or}\quad\frac{d}{dt}[K^{2}\|\dot{\mathbf{y}}\|^{2}]=0$
@@
In our numerical algorithms, we use Corollary 3.5 as an a-posteriori check to infer whether we have effectively solved (14); see Sections 4 and 5.

### v2

# Version of December 3, 2025

@@ Abstract.
In this work, we solve a discrete optimal transport problem in a nonuniform environment. To solve the optimal transport problem, we build the cost matrix and then use classical solvers for discrete optimal transport. The challenge is to form the cost matrix, which requires finding the optimal path between two points, and for this task we formulate and solve the associated Euler-Lagrange equations. A main contribution of ours is to provide verifiable sufficient conditions of optimality of the solution of the Euler-Lagrange equation and to propose new algorithms to to check optimality a-posteriori, thus validating the (exact) computation of the cost matrix. We illustrate our results and performance of the algorithms on several numerical examples in 2 and 3 dimensions.
@@
Optimal transport, numerical computation, optimal trajectory, calculus of variations, necessary and sufficient optimality condition, assignment problem, Sinkhorn method
@@ 1. Introduction
Motivated by solving optimal transport problems in the presence of obstacles, in this work we consider solving a discrete optimal transport of masses in a nonuniform environment. More specifically, we will consider the case where the cost function $c({\mathbf{a}},{\mathbf{b}})$ to move one unit of mass from location ${\mathbf{a}}\in{{\mathbb{R}}^{n}}$ to location ${\mathbf{b}}\in{{\mathbb{R}}^{n}}$ is given as the solution of a problem of calculus of variations of the form
(1) $c({\mathbf{a}},{\mathbf{b}})=\inf_{{\mathbf{x}}(t)\in{\mathcal{P}}({\mathbf{a}},{\mathbf{b}}}\int_{0}^{1}{\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})dt\ ,$
and we are interested in the special forms of $\mathcal{L}$} given in (5) and (6). From the theoretical point of view, the model (1) is a classical problem in Calculus of Variations (see Section 1.1 below), and it is considered also by Villani in his book on Optimal Transport, [VillaniOldNew, Chapter 7]. Our goal in this work is numerical, that is we want to compute the optimal transport plan of moving $k_{0}$ point masses ${\mathbf{x}}_{i}$’s into $k_{1}$ point masses ${\mathbf{y}}_{j}$’s, where ${\mathbf{x}}_{i}$’s and ${\mathbf{y}}_{j}$’s are points in some subset of ${\mathbb{R}}^{n}$, when the cost of moving one into the other is given by (1).
@@
In a uniform environment, the case that is usually considered in the literature, the transportation cost $c({\mathbf{x}},{\mathbf{y}})$ is typically given by a $p$-norm, $c({\mathbf{x}},{\mathbf{y}})=\|{\mathbf{x}}-{\mathbf{y}}\|_{p}$, $1<p<\infty$, the $2$-norm $\|{\mathbf{x}}-{\mathbf{y}}\|_{2}$, or simply $\|{\mathbf{x}}-{\mathbf{y}}\|$, being the most obvious choice, and also by the $2$-norm square, $c({\mathbf{x}},{\mathbf{y}})=\frac{1}{2}\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}$, a case that is known as the $2$-Wasserstein distance, or simply Wasserstein distance. So, in a uniform environment in ${\mathbb{R}}^{n}$}, computation of the cost is straightforward. But, in a nonuniform environment, the cost will generally depend on the path followed to go from $\mathbf{x}$} to $\mathbf{y}$}, and this makes the computation of $c({\mathbf{x}},{\mathbf{y}})$ a challenge in its own rights.
@@
Optimal transport, in one of its several variants, has been receiving a lot of attention in recent decades, in no small part because of its flexibility to adapt to many problems of seemingly different nature, such as the Schrödinger bridge problem, unbalanced optimal transport, transport in the presence of physical constraints, and the use of optimal transport in Machine Learning (e.g., see [PMLR-Neklyudov]). Our work is particularly motivated by studying optimal transport in the presence of physical constraints, obstacles. The study of optimal transport in the presence of obstacles has received both theoretical and computational attention; for example, see [Cavalletti] for theoretical results, and [BarrettPrigozhin, OT-convex-obstacle] for computational techniques. In particular, in [BarrettPrigozhin] Barrett and Prigozhin are concerned with transporting a mass into another, in the plane, by minimizing the cost given by the Euclidean distance when there is an obstacle along straight lines paths, and their computational approach consists of a regularization of the cost, followed by a finite element discretization of the underlying Monge-Ampere PDE. Presence of an obstacle is one instance of a nonuniform environment, which is the general case we consider, and we do it in the context of a discrete optimal transport problem so that we end up needing to compute optimal trajectories between two points (in the plane or in space) subject to a weighted “distance” function. This very task bears similarities with that of obtaining geodesics on general Riemannian manifolds (e.g., see [RiemannGeo, Section 2.C]), as we will exemplify in the body of the paper.
@@
A plan of the paper is as follows. In Section 1.1 below, we review the basic setup of the discrete optimal transport problem and discuss the two choices of cost we consider: the length and the energy costs. In Section 2, we derive the Euler-Lagrange equations for the cases of interest to us, recall a well known equivalence result between the two cases considered, and discuss its implications for algorithmic development. We also give an important theoretical result establishing sufficient conditions for optimality; we will exploit this result computationally to rigorously verify optimality. In Section 3, we outline the algorithms adopted in full details, and in Section 4 we present simulation results on several examples, highlighting performance on a number of scenarios.
1.1. Setup of the problem and our tasks
@@
1.1.1. Discrete Optimal Transport
This is the most commonly studied setup of an optimal transport problem.
@@ 1.1.1. Discrete Optimal Transport
Our numerical method will consist of forming the cost matrix $C=(c({\mathbf{x}}_{i},{\mathbf{y}}_{j}))_{i=1:k_{0},j=1:k_{1}}$ and then using well developed techniques to solve (4), namely an assignment solver or the Sinkhorn algorithm, see Section 3.3. So doing, we need to effectively form and store the cost matrix. Of course, storing the matrix $C$ may not be feasible if $k_{0}$ and $k_{1}$ are truly large; but, for the values of $k_{0}$ and $k_{1}$ we have in mind (say, in the order of thousands) this is not a concern. As far as forming $C$, this will be discussed in details below. Presently, we just observe that forming $C$ is a very parallelizable task, since the entries $c({\mathbf{x}}_{i},{\mathbf{y}}_{j})$ are independent of each other. So, the cost of computing $C$ is proportional to that of computing one trajectory (assuming that computation of trajectories costs roughly the same) times $k_{0}k_{1}$ and divided by the number of processors we have. Moreover, we want to form $C$ “exactly” in nonuniform environments, and this will require carefully validated computation of the $c({\mathbf{x}}_{i},{\mathbf{y}}_{j})$’s. Of course, there are alternatives to our approach that avoid forming $C$ directly. For example, the authors of [SolomonEtAl] – who are concerned with computer graphics applications and large values of $k_{0}$ and $k_{1}$ – develop a method that considers regularized Wasserstein distances by computing optimal transportation plans through iterative kernel convolutions, by adapting the approach of [BB]; they give impressive graphical displays of their results, but their work lacks a convergence analysis and it does not appear to be immediately applicable to nonuniform environments.
1.1.2. Cost Matrix
To set up the cost matrix, the first thing to do is to select a cost $c({\mathbf{a}},{\mathbf{b}})$. We consider two costs, the “length”, and the “energy” costs, which generalize the Euclidean and the Wasserstein distances, respectively.
Agreement. In this work, the functional $K$ below is as smooth as required to validate differentiating it as often as needed.
The length cost. This is given by
@@ 1.1.2. Cost Matrix
where the norm is the Euclidean norm, and $K$ is a positive function characterizing the environment, that is its geometry. If $K=1$ (or any fixed constant, for what matters), we have a uniform environment, otherwise a nonuniform environment. Clearly, if $K=1$, we are just looking at the arc-length of the path from $\mathbf{a}$} to $\mathbf{b}$}, whereas if $K$ is not constant we have a weighted arc-length of the path. It is a simple exercise in calculus of variations (e.g., see [Liberzon]) that, when $K=1$, the shortest path is a straight line: ${\mathbf{x}}(t)={\mathbf{a}}+t({\mathbf{b}}-{\mathbf{a}})$, $0\leq t\leq 1$, so that the cost is precisely the Euclidean distance from $\mathbf{a}$} to $\mathbf{b}$}: $c({\mathbf{a}},{\mathbf{b}})=\|{\mathbf{a}}-{\mathbf{b}}\|$, as one should expect in a uniform environment.
The energy cost This is given by
@@ 2. Optimal Trajectory
Our main concern in this section is to solve the minimization problem for the optimal trajectory from $\mathbf{a}$} to $\mathbf{b}$} for $\mathbf{x}$}-dependent weight function $K({\mathbf{x}})$, which we need to do in order to setup the cost matrix. With respect to (1), ${\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})$ is either the integrand in (5) or (6). Our task is to find the trajectory ${\mathbf{x}}\in{\mathcal{P}}({\mathbf{a}},{\mathbf{b}})$ which gives the minimum value below
@@
Regardless of whether $\mathcal{L}$} is given by (5) or by (6), the minimizing path will be called a geodesic; this naming will be justified by Fact 2.3 below.
@@ 2.1. Necessary condition: Euler-Lagrange equations
The problem formulated in (7) is a classical problem of Calculus of Variations and the search for a minimizer follows the approach based on solving the Euler-Lagrange equations, which are first order necessary conditions for obtaining a minimum. Naturally, the derivation of the Euler-Lagrange equations below is a standard textbook procedure, but since we have not explicitly found the precise form of the equations for the Lagrangian $\mathcal{L}$} given by (5) or by (6), and their derivation is short, we provide details. For (7), the Euler-Lagrange equations are:
(8) $\frac{\partial{\mathcal{L}}}{\partial x_{i}}-\frac{d}{dt}\bigg[\frac{\partial{\mathcal{L}}}{\partial\dot{x}_{i}}\bigg]=0\ ,\ i=1,2,\dots,n\ ,$
or $\sum_{j=1}^{n}{\mathcal{L}}_{\dot{x}_{i}\dot{x}_{j}}\ddot{x}_{j}+\sum_{j=1}^{n}{\mathcal{L}}_{\dot{x}_{i}x_{j}}\dot{x}_{j}-{\mathcal{L}}_{x_{i}}=0$, for all $i$, or more compactly as
@@
where $D^{2}_{\dot{\mathbf{x}}}{\mathcal{L}}\in{{\mathbb{R}}^{n\times n}}$, $(D^{2}_{\dot{\mathbf{x}}}{\mathcal{L}})_{ij}={\mathcal{L}}_{\dot{x}_{i}\dot{x}_{j}}$, and $D_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathcal{L}}\in\mathbb{R}^{n\times n}$, $(D_{\dot{\mathbf{x}}{\mathbf{x}}}{\mathcal{L}})_{ij}={\mathcal{L}}_{\dot{x}_{i}x_{j}}$. We can further simplify (9) in the two cases of interest to us.
@@ 2.1.2. Euler-Lagrange for energy cost
Here
(13) ${\mathcal{L}}({\mathbf{x}},\dot{\mathbf{x}})=\frac{1}{2}K^{2}({\mathbf{x}})\|\dot{\mathbf{x}}\|^{2}\ ,$
and it is easier to rewrite the boundary value problem (9).
@@
$K^{2}({\mathbf{x}})\ddot{\mathbf{x}}+2K({\mathbf{x}})\dot{\mathbf{x}}[\nabla K]^{T}\dot{\mathbf{x}}=K({\mathbf{x}})\|\dot{\mathbf{x}}\|^{2}\nabla K\ ,$
@@
$\ddot{\mathbf{x}}=\frac{\|\dot{\mathbf{x}}\|^{2}}{K({\mathbf{x}})}\nabla K-\frac{2(\nabla K)^{T}\dot{\mathbf{x}}}{K({\mathbf{x}})}\dot{\mathbf{x}}\ ,$
@@
An interesting and useful fact about solutions of (14) is that they have constant speed.
@@
$\ddot{\mathbf{y}}=\frac{\|\dot{\mathbf{y}}\|^{2}}{K({\mathbf{y}})}\nabla K({\mathbf{y}})-\frac{2[\nabla K({\mathbf{y}})]^{T}\dot{\mathbf{y}}}{K({\mathbf{y}})}\dot{\mathbf{y}}\ ,$
@@
$K\ddot{\mathbf{y}}^{T}\dot{\mathbf{y}}=-\|\dot{\mathbf{y}}\|^{2}[\nabla K]^{T}\dot{\mathbf{y}}\quad\text{or}\quad\frac{d}{dt}[K^{2}\|\dot{\mathbf{y}}\|^{2}]=0\ ,$
@@
In our numerical algorithms, we use Corollary 2.5 as an a-posteriori check to infer whether we have effectively solved (14); see Sections 3 and 4.